OpenStudy (anonymous):

Find the 21st term of each arithmetic sequence -1.45 , -2.9 , -4.35 , -5.8 , -7.25

hartnn (hartnn):

do you see any pattern in this sequence ?

OpenStudy (anonymous):

what is the common difference

OpenStudy (anonymous):

nope , i have like 5 of these and i did 3 of them and i cant do the rest , including this one , i dont get this one.

hartnn (hartnn):

know what common difference is ??

hartnn (hartnn):

common difference(d) is the difference between consecutive term d= 2nd term - 1st term = 3rd term - 2nd term = ....and so on

hartnn (hartnn):

here, d = -2.9- (-1.45) = ...

OpenStudy (anonymous):

= -1.45 ?

hartnn (hartnn):

yes, correct d = -1.45 now just use the general formula , \(\Large a_n = a_1 +(n-1)d\) a1 = 1st term = -1.45 n = 21 d= -1.45

OpenStudy (anonymous):

oh i got it , thanks.!!!!

hartnn (hartnn):

welcome ^_^